Circular No. 8744

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 R1 (SIDING SPRING) R. H. McNaught reports the discovery of a comet (discovery observa-

tion tabulated below) on images taken by D. M. Burton with the 0.5-m Uppsala Schmidt telescope in the course of the Siding Spring Survey, the object reported as slightly diffuse with an apparent 10′′ coma. Additional images taken on Sept. 2.55 UT show the object to show a consistent ‘softness’ and be noticeably larger than stars of similar brightness. McNaught adds that an image taken by Burton on Sept. 2.725 shows the object to be “barely, but certainly, diffuse in images taken after moonset”.

2006 UT  α_{2000}  δ_{2000}  Mag.
Sept.1.49494  22^{h}51^{m}49.^{s}46  −53^{o}09′50′′.0  17.7

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2006-R08.

T = 2006 Sept. 3.244 TT  ω = 251.^{o}186
Ω = 220.610 }2000.0
q = 1.69969 AU  i = 159.844

SUPERNOVAE 2006ep, 2006eq, 2006er Several new apparent supernovae have been reported: 2006ep indepen-

dently on unfiltered CCD images by N. Joubert and W. Li (LOSS/KAIT; cf. IAUC 8741; tabulated data below from KAIT) and by K. Itagaki (cf. IAUC 8711; via S. Nakano, Sumoto, Japan); and 2006eq and 2006er on g-, r-, and i-band images by J. Frieman (on behalf of the Sloan Digital Sky Survey II collaboration; details given on CBET 611; approximate g magnitudes tabulated below).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006ep Aug. 30.35 0^{h}41^{m}24.^{s}88 +25^{o}29′46′′.7 17.8 43′′ W, 11′′.3 S
2006eq Aug. 26  21 28 37.13 + 1 13 41.2 18.7 7′′ W, 7′′ S
2006er Aug. 26  0 21 37.54 − 1 00 35.9 19.5 5′′ W, 2′′ N

Additional magnitudes for 2006ep in NGC 214: Aug. 21.6 UT, [19 (Itagaki); 25.37, [19.2 (KAIT); 26.64, [19 (Itagaki); 30.637, 18.1 (Itagaki); 31.60, 17.8 (Itagaki). Itagaki’s position end figures for 2006ep are 24^{s}.88, 46′′.5. SN 2006ep is a type-Ib/c supernova (cf. CBET 612, 614), now near maximum. SN 2006eq is of type Ia, now within a week after maximum (cf. CBET 612).

2006 September 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green